### v2

@@ 6.5.1 DEFINITION
Cross-domain AI, also known as multi-domain AI, is defined as the integration of collaborative AI-enabled functionalities across multiple domains [oran-multidomain]. In wireless networks, this concept aims to jointly optimize network performance across RAN, TPN, and CN, as shown in Fig. 6. Cross-domain AI involves the integration of information, data exchange, collaborative model training and inference, as well as the sharing of computing resources. It is noted that cross-domain AI includes tasks such as power allocation in RAN, routing strategies in TPN, and slicing configurations in CN. By enabling coordinated decision-making across these domains, cross-domain AI can contribute to seamless multi-vendor integration and large-scale network optimization, thereby supporting scalable and resource-efficient ORAN infrastructures.
6.5.2 RECENT ADVANCES AND PROMISING DIRECTIONS
@@
The proposed cross-domain SMO layer in Fig. 6 inherently enables cross-domain AI capabilities. However, several critical challenges remain before fully exploiting the advantages of cross-domain AI. First, each domain holds distinct optimization objectives. For instance, the RAN domain focuses on reliability, latency, and throughput, while the CN focuses on meeting service-level agreements (SLAs) for emerging applications such as AR/VR and real-time automation. Hence, designing an E2E optimization framework that balances these domain-specific objectives remains an important direction for future research. Second, the cross-domain AI controller must process heterogeneous data flows collected at different timescales. This necessitates the design of customized AI models that incorporate domain-specific knowledge in each data flow. Meanwhile, future work should also investigate the data flow from TPN and CN to achieve efficient fusion processing.
@@
The open architectures, cloud-native deployments, and cross-domain data sharing for AI orchestration may also introduce new vulnerabilities that compromise network security and privacy [oran-wg11]. Major mobile network operators are adopting zero trust architecture (ZTA) principles in their ORAN deployments. The O-RAN Alliance’s Security Working Group (WG11) mandates ZTA principles for all compliant components, enforcing mutually-authenticated transport layer security (mTLS), least-privilege access control, and standardized encryption, such as TLS 1.2/1.3 and SSH v2, to ensure secure data exchange across open interfaces. These measures help mitigate risks in the multivendor and disaggregated environments [oran-wg11]. Emphasis on the secure software development lifecycle (SDLC), and the incorporation of software bill of materials (SBOMs) for third-party rApps and xApps has also become a central focus for vendors and operators in their procurement and vetting processes. To protect privacy, operators must anonymize or aggregate sensitive network data before sharing it with external applications. The onboarding and validation of third-party rApps and xApps in practice include containerization, sandboxing, and runtime monitoring to prevent any malicious application from compromising the network. This forms a vital security layer for any operator running multivendor apps on the RIC [nist-sbom]. Meanwhile, federated learning is being explored to optimize spectrum allocation and resource management without centralizing sensitive user data by major operators and vendors. However, the commercial-scale adoption of federated learning for security and large-scale network optimization remains in its early stages, requiring further research in future work [fl-privacy-2025].
6.6 DIGITAL TWIN
6.6.1 DEFINITION
@@
The digital twin (DT) is a virtual replica of a physical system that enables simulation, analysis, and optimization of system behavior in a risk-free environment [tran2025network]. Specifically, telecom companies often evaluate new network setups in controlled laboratory environments with limited data and a small number of users. This approach, commonly referred to as a pilot environment, enables service providers to assess the interaction of new technologies with existing systems, without risking large-scale customer disruptions. Although pilot testing has inherent limitations, direct testing in operational networks carries more substantial risks. The failures during direct testing may lead to service outages and significant costs, including revenue loss, customer compensation, and reputational damage. Therefore, implementing DT in wireless networks has become essential for operators and users, since it can help offer a balance between the realism of pilot environments and the risk mitigation required for large-scale deployment.
@@
With a DT-based ORAN for NTNs, operators can implement different regenerative architectures and evaluate end-to-end performance under diverse wireless channels. DTs further empower ORAN to perform network analytics and diagnosis for high-efficiency service in highly dynamic NTN scenarios [zhou2025digital]. By leveraging the ORAN functional split, the physical NTN can be flexibly represented in its DT-NTN, which can enhance the accuracy of simulations and the capabilities of predictive and optimization [Hraishawi2025].
6.6.2 RECENT ADVANCES AND PROMISING DIRECTIONS
@@
DTs are envisioned as a key enabler for intelligent network design, analysis, operation, and automation, complementing AI-native management frameworks in future 6G networks. Recent research has highlighted the growing importance of AI, DT, and ORAN for 6G NTN scenarios [campoy2025digital, wu2024ai, tao2024wireless, moore2025next, mushi2024open, Hraishawi2025, firouzi20252]. Specifically, several studies have explored AI-driven approaches to enhance DT accuracy and scalability. For instance, reinforcement learning (RL) has been integrated into scalable 6G RAN DTs to enable adaptive decision-making in [campoy2025digital], and the authors in [wu2024ai] highlighted the integration of ML, deep learning (DL), and natural language processing (NLP) in DTs. Generative AI (GenAI) models such as transformers and diffusion architectures have been employed to construct hierarchical DT frameworks in [tao2024wireless]. In addition, DT-based emulation platforms and testbeds have been developed to support ORAN-based NTN experimentation in [moore2025next, mushi2024open]. The DT-enabled ORAN framework has been investigated in terms of its effectiveness in optimizing resource allocation [Hraishawi2025] and facilitating mission-aware autonomy in dynamic NTN environments [firouzi20252].
@@
As discussed above, ORAN enables on-orbit tests and upgrades for NTN and thus avoids the costly and time-consuming launch process. However, on-orbit tests and upgrades in real networks still face much higher risks than traditional terrestrial networks due to large-scale coverage and complex environments, thereby necessitating high-fidelity DT techniques for NTNs [liu2024challenges]. It is noted that the Near-RT RIC inherently facilitates the creation of a virtual replica of the ORAN-based NTN by providing AI inference capability. To guarantee the accuracy of inference, the Near-RT RIC is required to achieve stringent synchronization with the physical systems, including O-RU, O-DU, and O-CU. Meanwhile, unlike the traditional terrestrial DT with adequate communication and computation capabilities in the central server, the Near-RT RIC must guarantee the execution of xAPPs and the existing communication overheads with other components. Apart from that, the mobility of Near-RT RIC also leads to the synchronization challenge along with the mobility of O-RU, O-DU, and O-CU [Hraishawi2025]. To solve the challenges above, a hierarchical DT framework should be developed for ORAN-based NTN. Specifically, the DT of Near-RT RIC can be constructed first to predict its mobility and available communication/computation resources. Then, the Near-RT RIC schedules the communication and computation resources to construct the DT of ORAN-based NTN for inference.
7 CONCLUSIONS
@@
The development of efficient and flexible NTN architectures is essential for future AI-based wireless networks to achieve ubiquitous coverage and support advanced use cases. This survey has presented a comprehensive and structured overview related to the development of the AI-Native ORAN-based NTN framework, supporting dynamic configuration, scalability, and intelligent orchestration in future 6G NTN systems. Specifically, this survey first reviewed the state-of-the-art research on ORAN-based NTNs, highlighting their design principles, technical progress, and existing limitations. It also offered necessary background knowledge and key related literature on ORAN, NTN, and AI-Native for communications. Furthermore, this paper analyzed the unique challenges encountered across the DevOps lifecycle of NTNs and proposed an orchestrated AI-Native ORAN-based NTN framework. Meanwhile, the key enabling technologies of the proposed framework, including dynamic wireless FH splits, multi-tier RICs for distributed learning, flexible functional role switch, and cross-domain SMO, were discussed in detail. Finally, this paper presented several representative use cases to demonstrate the applicability of the proposed framework, and outlined prospective future research directions to further enhance the AI-Native network control, operation, and optimization.
@@ ACKNOWLEDGMENT
The authors would like to thank the support of the interdisciplinary research center for communication systems and sensing (IRC-CSS), King Fahd University of Petroleum and Minerals. The authors would also like to thank the support of the Office of Sponsored Research, King Abdullah University of Science and Technology.

### v3

@@ 6-D.1 DEFINITION
Cross-domain AI, also known as multi-domain AI, is defined as the integration of collaborative AI-enabled functionalities across multiple domains [94]. In wireless networks, this concept aims to jointly optimize network performance across RAN, TPN, and CN. Cross-domain AI involves the integration of information, data exchange, collaborative model training and inference, as well as the sharing of computing resources. It also includes tasks such as power allocation in RAN, routing strategies in TPN, and slicing configurations in CN. By enabling coordinated decision-making across these domains, cross-domain AI can contribute to seamless multi-vendor integration and large-scale network optimization, thereby supporting scalable and resource-efficient ORAN infrastructures.
6-D.2 RECENT ADVANCES AND PROMISING DIRECTIONS
@@
The proposed cross-domain SMO layer in Fig. 6 inherently enables cross-domain AI capabilities. However, several challenges remain before fully exploiting the advantages of cross-domain AI. First, each domain has different optimization objectives. For instance, the RAN domain focuses on reliability and throughput, while the CN focuses on meeting SLAs for emerging applications such as AR/VR and real-time automation. Hence, designing an E2E optimization framework that balances these domain-specific objectives remains an important research direction. Second, the cross-domain AI controller must process heterogeneous data flows collected at different timescales. This necessitates the design of customized AI models that incorporate domain-specific knowledge in each data flow.
@@
Moreover, in cross-domain AI, the open architectures and cross-domain data sharing may also introduce new vulnerabilities that compromise network security and privacy. To address these risks, several security mechanisms have been established. From a standardization perspective, the O-RAN Alliance’s Security Working Group (WG11) mandates zero trust architecture (ZTA) principles for all compliant components, enforcing mutually-authenticated transport layer security (mTLS), least-privilege access control, and standardized encryption, such as TLS 1.2/1.3 and SSH v2, to ensure secure data exchange across open interfaces [88]. From the perspective of vendors or operators, they emphasize the secure software development lifecycle (SDLC) and the use of software bills of materials (SBOMs) when procuring and vetting third-party rApps and xApps [86]. To protect privacy, sensitive network data must be anonymized or aggregated by operators before being shared with external applications. In practice, the onboarding and validation process of these rApps and xApps would require containerization, sandboxing, and runtime monitoring to prevent any malicious application from compromising the network. These mechanisms collectively form a key security layer for operators running multivendor applications on the RIC.
Furthermore, AI models embedded in RICs, xApps, or rApps may also face adversarial attacks or model failures. To enhance the AI robustness or safety, future research can focus on advancing model verification techniques and resilient inference strategies. In addition, federated learning is being explored by major operators and vendors to optimize resource management in a privacy-preserving manner, without requiring the centralization of sensitive user data [66]. However, its commercial-scale adoption for secure and large-scale network optimization is still at an early stage, highlighting the need for further research.
6-E MOBILITY MANAGEMENT
6-E.1 DEFINITION
Mobility management is a critical challenge in NTNs, mainly encompassing handover, location management, and mobility robustness optimization. In TN, handover decisions are primarily coverage-driven and are typically triggered by the degradation of signal strength indicators such as RSRP and RSRQ. As the user equipment (UE) approaches the cell edge, the serving BS relies on UE measurement reports and predefined handover strategies to determine whether a handover should be initiated [53].
However, in the satellite-based NTN, mobility is primarily driven by satellite beam movement rather than the UE’s motion. For LEO satellites orbiting at 7.8 km/s, beams sweep rapidly across the Earth’s surface, causing even stationary users to experience frequent handovers every few minutes. This high handover frequency, together with large Doppler shifts, leads to increased UE power consumption and considerable signaling overhead [17]. In addition, the relatively constant UE-satellite distance within a beam footprint leads to minimal signal variations, making conventional measurement-based handover criteria less effective. Furthermore, inherent NTN characteristics such as long propagation delays, dynamic beam patterns, and intermittent feeder links may lead to outdated measurements and handover failures.
6-E.2 RECENT ADVANCES AND PROMISING DIRECTIONS
Recent studies have focused on improving mobility management in NTNs to ensure reliable connectivity [17, 109, 97, 116]. Considering the different payload deployment options in satellite-based NTN and the ORAN architecture, the handover procedures have been categorized into intra-DU (intra-satellite), inter-DU (inter-satellite), and inter-CU (inter-access) handovers in [97]. 3GPP has specified several NTN mobility management trigger mechanisms, based on measurement, location, time, timing advance, and elevation angle, respectively [116]. In [17], the authors discussed different mobility management solutions for satellite-based 5G NTNs and evaluated their impacts on UE power consumption and signaling overhead, highlighting the potential of AI-based methods for dynamic mobility management optimization. The authors in [109] focused on optimizing handover performance in NTN to enhance UE effective service time, and also investigated the impact of different ORAN functional splits between ground stations and LEO satellites on handover delays. This work revealed that the highest availability was achieved with the gNB onboard the satellite.
The above studies have advanced mobility management in NTNs from different perspectives. To achieve more effective mobility management within the proposed ORAN-based NTN framework, several research directions can be further explored. First, ephemeris-aware xApps should be developed to support proactive beam and satellite handovers, and Near-RT RIC policies need to be designed to mitigate Doppler and propagation delay effects [24]. Second, AI/ML models can be leveraged to explore the multi-constellation mobility mechanisms for seamless connectivity. In addition, lightweight inference frameworks are also required to ensure reliable decision-making under the stringent backhaul constraints.
6-F DIGITAL TWIN
6-F.1 DEFINITION
@@
The digital twin (DT) is a virtual replica of a physical system that enables simulation, analysis, and optimization of system behavior in a risk-free environment [120]. For telecom companies, the direct testing of new network setups in operational networks usually involves high risks, since its failures may lead to service outages, significant revenue loss, and reputational damage. As a result, new network configurations are often evaluated in controlled laboratory environments with limited data and a small number of users. This approach, commonly referred to as a pilot environment, enables service providers to assess the interaction of new technologies with existing systems, without risking large-scale customer disruptions. To provide realistic system evaluation while further mitigating operational risks, implementing DT in wireless networks has become essential for operators and users.
@@
In the proposed ORAN-based NTN framework, the integration of DTs helps operators implement different regenerative architectures and evaluate E2E performance under diverse wireless channels. In addition, DTs also further empower ORAN to perform network analytics and diagnosis for high-efficiency service in highly dynamic NTN scenarios [135].
6-F.2 RECENT ADVANCES AND PROMISING DIRECTIONS
@@
DTs are envisioned as a key enabler for intelligent network design, analysis, operation, and automation. They complement the AI-Native management frameworks in future 6G networks. Recent research has highlighted the growing importance of AI, DT, and ORAN for 6G NTN scenarios [27, 130, 119, 82, 85, 12, 47]. Specifically, several studies have explored AI-driven approaches to enhance DT accuracy and scalability. For instance, reinforcement learning (RL) has been integrated into scalable 6G RAN DTs to enable adaptive decision-making in [27], and the authors in [130] highlighted the integration of ML, deep learning (DL), and natural language processing (NLP) in DTs. Generative AI (GenAI) models such as transformers and diffusion architectures have been employed to construct hierarchical DT frameworks in [119]. In addition, DT-based emulation platforms and testbeds have been developed to support ORAN-based NTN experimentation in [82, 85]. The DT-enabled ORAN framework has been investigated in terms of its effectiveness in optimizing resource allocation [12] and facilitating mission-aware autonomy in dynamic NTN environments [47].
@@
As discussed above, ORAN enables on-orbit tests and upgrades for NTN and thus avoids the time-consuming launch process. However, on-orbit tests and upgrades in the real and complicated NTN environment face much higher risks than traditional TN, which highlights the need for high-fidelity DT technologies [76]. Within the proposed framework, the Near-RT RIC has the AI inference capability to create a virtual replica of an ORAN-based NTN, but also faces several challenges. Specifically, to guarantee the inference accuracy, the Near-RT RIC needs to maintain strict synchronization with the physical network elements O-RU/O-DU/O-CU under this highly mobile scenario. In addition, unlike the traditional terrestrial DT with adequate communication and computation capabilities in the central server, the Near-RT RIC in NTNs must also guarantee the xApps execution and the existing communication overhead with other network components. To address the challenges above, a hierarchical DT framework should be developed for ORAN-based NTN. To be more specific, a DT of Near-RT RIC is first established to model and predict its mobility as well as the available communication and computation resources. Based on these predictions, the Near-RT RIC subsequently schedules the available resources to construct a system-level DT of the ORAN-based NTN for inference tasks.
7 CONCLUSION
@@
Developing efficient and flexible NTN architectures is essential for AI-based wireless networks to achieve ubiquitous coverage and support advanced 6G use cases. This survey has presented a comprehensive and structured overview related to the development of the AI-Native ORAN-based NTN framework, supporting dynamic configuration, scalability, and intelligent orchestration in future 6G NTN systems. Specifically, this survey first reviews the state-of-the-art research on ORAN-based NTNs, highlighting their design principles, technical progress, and existing limitations. It also offers necessary background knowledge and key related literature on ORAN, NTN, and AI-Native for communications. Furthermore, this paper analyzes the unique challenges encountered across the DevOps lifecycle of NTNs and proposes an orchestrated AI-Native ORAN-based NTN framework. Meanwhile, the architecture and key technological enablers of the proposed framework, including dynamic wireless FH splits, multi-tier RIC structure, flexible functional role switch, and cross-domain SMO, are discussed in detail. More importantly, this paper presents several representative use cases to demonstrate the applicability of the proposed framework. Finally, this paper outlines prospective future research directions to further enhance the AI-Native network control, operation, and optimization in NTN.
@@ ACKNOWLEDGMENT
The authors would like to thank the support of the interdisciplinary research center for communication systems and sensing (IRC-CSS), King Fahd University of Petroleum and Minerals (KFUPM). The authors would also like to thank the support of the Office of Sponsored Research, King Abdullah University of Science and Technology (KAUST).